if a and b are integers and 14/3=a/b, which of the following could be value of a/b?
A.68
B.66
c.63
D.60
E.58

Answer by jim_thompson5910(35256)

If 14/3 = a/b or a/b = 14/3, then clearly a/b is NOT an integer. So a/b cannot be any of the answer choices shown below. There must be a typo somewhere.